# Canonical Electrodynamics in Ashtekar-Barbero variables

### v1

@@ Abstract
We present the full Hamiltonian analysis of electrodynamics, including fermions, photons, and their interaction, on a general curved spacetime in Ashtekar-Barbero variables. No gauge fixing is performed and the theory is background independent. The details of the constraints of the theory for both Weyl and Dirac fermions are presented. Moreover, the equations of motion are derived and it is shown that they reduce to the equations of motion of the standard QFT in curved or flat spacetime in the appropriate limits. Finally, the behavior of the theory under parity transformation is analyzed.
@@ 1 Introduction
There are previous works done with regard to fermions in a canonical theory in Ashtekar-Barbero variables Bojowald:2010qpa ; Das:2008r ; Thiemann:1998QSDV ; Jacobson:1988qta ; Mercuri:2006wb ; PhysRevD.73.084016 ; Date:2011rd that paved the way for our work. However, they usually ignore the interaction term between the fermions and the photons and hence do not present the full QED Hamiltonian in a non-fixed general curved background. Some of them also lack the details of computation or seemingly suffer from some missing terms or some inconsistencies. In particular, it seems that in Das:2008r the choice of the definition of the spin connection $\Gamma_{a}^{i}$ is not compatible with their covariant derivative $D_{a}=\partial_{a}+A_{a}^{i}\tau_{i}$ and their constraints. In Thiemann:1998QSDV , the authors present the Hamiltonian constraint for gravity plus fermions, however, it seems that they do not start directly from the action, and as a consequence there is a missing term in the modification of the Ashtekar connection as well as in the fermionic Hamiltonian constraint. They also seem to take the modification of the Ashtekar connection derived in Jacobson:1988qta . However, the latter starts from the Palatini action rather than the Holst action . As a consequence it does not seem to involve axial fermionic current and only incoporates fermionic current. It also seems that the fact that the extrinsic curvature in presence of fermions is no longer symmetric because of the torsion contribution is overlooked. As a consequence, in the scalar constraint, a term proportional to $\epsilon_{i}{}^{jk}K_{a}^{i}E_{j}^{a}J_{k}$ is missing.
@@
In this work we set to fill the aforementioned gaps. We carefully study a generic background-independent system of gravity, fermions and photons, starting from the action and making a $3+1$ decomposition, writing the theory in Ashtekar-Barbero variables. This is done for both Weyl and Dirac fermions, in a step by step, clear, and detailed manner. We then derive the modification to both the canonical variables, and the constraints, showing in a clear way the distinct contribution of each sector of to these constraints. We also provide the equations of motion of fermions in this setting and study the behavior of the system under parity transformation. Another difference of this work compared to the previous ones, is that we do not modify the gamma matrices, and rather change the identities between them to incorporate the changes needed for the background-independent framework. This way it is easier to go back to a flat QFT theory. Furthermore, as mentioned before, to our knowledge, this is the first time that the QED interaction term is studied in detail in such a canonical formulation in Ashtekar-Barbero variables.
@@
The structure of this paper is as follows: In Sec. 2, we provide a brief overview of vacuum canonical gravity in Ashtekar-Barbero variables starting from the Holst action. Then in Sec. 3.1, we present how Weyl fermions are incorporated in the theory, and how they modify the canonical variables and constraints. In Sec. 3.2 a similar analysis is provided but now for Dirac fermions. Using the results of the previous sections, in Sec. 4.1 the full QED Hamiltonian in an arbitrary dynamical background is constructed by incorporating both the photon field and its interaction with fermions and gravity. In Sec. 4.2 we study the behavior of the system under the parity transformations. In Sec. 4.3 we briefly discuss how the constraints of the gravity-fermion system would modify in case of non-minimal coupling. Finally, in Sec. 5 we summarize and conclude the paper.
@@ 2 Brief overview of Ashtekar-Barbero variables
In this section we will quickly review Ashtekar-Barbero formalism of GR in order to introduce the notation used in the Fermions and QED Sections.
We start with a first order action of GR, the Holst action Bojowald:2010qpa
@@
Throughout this work the first four lower case Latin indices $a,\,b,\,c,\,d$ are abstract spacetime (sometimes spatial) indices. The rest of the lower case Latin indices are $su(2)$ indices. Upper case Latin indices are the internal Lorentz indices associated to the $SO(1,3)$ gauge freedom under which the tetrads transform. The connection one form associated to this transformation is $\omega_{a}^{IJ}$, and $F^{IJ}(\omega)=d\omega^{IJ}+\omega^{I}{}_{K}\wedge\omega^{KJ}$ is the antisymmetric 2-curvature of $\omega_{a}^{IJ}$. Furthermore $\kappa=8\pi G$ and
@@
This condition expresses the compatibility between the tetrad and the spin connection. This ensures that all tensor fields (such as the Dirac spinor field, the metric, etc.) transform properly under local Lorentz transformations.
To proceed toward a canonical formulation, one foliates the spacetime into spatial slices following the ADM decomposition. For this purpose, a spacetime tensor field, the triad $\mathcal{E}_{I}^{a}=e_{I}^{a}+n_{I}n^{a}$, is introduced where $n^{a}$ is the unit timelike vector field normal to the spatial slices, and $n_{I}\coloneqq e_{I}^{a}n_{a}$. It is also common to fix the boost part of the internal Lorentz transformations by imposing the “time gauge” condition $e_{0}^{a}=n^{a}$ Bojowald:2010qpa .
@@
where $N$ is the lapse function and $N^{a}$ the shift vector. Making use of the relation $|\text{det}\,e|=N\sqrt{\text{det}\,h}$, where $\text{det}\,h$ is the determinant of the spatial metric $h_{ab}$ induced on the spatial hypersurfaces, the action becomes
@@
3 Fermions
3.1 Weyl Fermions
We now study how the inclusion of Weyl fermions alters the familiar vacuum scenario. This system has been previously studied in the literature Bojowald:2010qpa ; Das:2008r ; Thiemann:1998QSDV ; Jacobson:1988qta ; Mercuri:2006wb ; PhysRevD.73.084016 ; Date:2011rd . Notably, in Bojowald:2010qpa and Das:2008r , the Dirac matrices themselves were modified to accommodate the curved spacetime geometry. In contrast, the present work adopts an approach in which instead of altering the matrices, we modify the algebraic identities involving them (see Appendix A). This results in a formulation that more closely mirrors the flat spacetime treatment of fermionic fields, allowing a more straightforward comparison between the two cases.
The starting point is the study of the Dirac action discussed in Appendix A.3, specifically Eq.(183). In order to derive the constraints associated to the fermionic contribution, we must revisit the canonical decomposition $\mathcal{E}_{I}^{a}=e_{I}^{a}+n^{a}n_{I}$ where $n_{I}=-\delta_{I}^{0}$ (in the time gauge), as was done in the vacuum case. As a result, the fermionic action becomes
@@ 3.1 Weyl Fermions
Once again we encounter three types of terms involving $t^{a},\,N$, and $N^{a}$ The term containing $t^{a}$ leads to
@@
$\mathcal{H}^{F}=$ $-\frac{i}{2}E_{i}^{a}\left(\overline{\Psi}\gamma^{i}\mathfrak{D}_{a}\Psi-\overline{\mathfrak{D}_{a}\Psi}\gamma^{i}\Psi\right)\,,$ (30)
@@
By minimally coupling the vacuum action (see Section 2) to the fermionic action, we obtain the total Hamiltonian of the system, comprising both the vacuum and the (Weyl) fermionic field contributions.
@@
$-\frac{i}{2}E_{i}^{a}\left(\overline{\Psi}\gamma^{i}\mathfrak{D}_{a}\Psi-\overline{\mathfrak{D}_{a}\Psi}\gamma^{i}\Psi\right),$ (35)
@@
For the sake of notational simplicity, we will omit the superscript “G+F” in the remaining of this section. Unless otherwise stated, all expressions should be understood as referring to the total system, including both the vacuum and the (Weyl) fermionic fields.
@@
This implies that the system is no longer torsionless and the canonical variables of the theory now carry an implicit dependence on the fermionic field. Moreover, as already noted in Perez:2005pm , unlike the vacuum case, where the Barbero-Immirzi parameter plays no role in the equations of motion, here the torsion term is proportional to this parameter, effectively governing the strength of the interaction between the gravitational and fermionic fields. As a result, to enforce the compatibility condition in the presence of fermions, the connection 1-form must be modified as
@@
$=\beta K_{a}^{i}+\tilde{\Gamma}_{a}^{i}+\frac{\kappa}{4}\frac{\beta^{2}}{1+\gamma^{2}}\left(e_{a}^{i}J^{0}-\frac{1}{\beta}\epsilon^{i}{}_{jk}e_{a}^{j}J^{k}\right)\,.$
@@
Once again, we see that the $\mathcal{S}_{i}$ constraint (34) is proportional to the Gauss constraint and can be neglected. Finally, the first class [^2] A constraint is said to be first class if its Poisson bracket with all other constraints vanishes weakly. constraints in the case in which gravity is minimally coupled to a fermionic field are
@@
$\mathcal{H}=$ $\frac{1}{2\kappa}\frac{E_{i}^{a}E_{j}^{b}}{\sqrt{\text{det}\,h}}\epsilon^{ij}{}_{k}\left(\mathcal{F}_{ab}^{k}(A)-(1+\beta^{2})\epsilon^{k}{}_{mn}K_{a}^{m}K_{b}^{n}-2\frac{1+\beta^{2}}{\beta}\mathcal{D}_{[a}^{(\Gamma)}K_{b]}^{k}\right)$
$-\frac{i}{2}E_{i}^{a}\left(\overline{\Psi}\gamma^{i}\mathfrak{D}_{a}\Psi-\overline{\mathfrak{D}_{a}\Psi}\gamma^{i}\Psi\right)\,,$ (45)
@@
Another modification arises in the definition of the fermionic covariant derivatives, which can now be reexpressed in terms of the covariant derivative associated with the modified Ashtekar connection, along with an additional correction term
@@
Accordingly, for the conjugate fermionic field, we find
@@
$\mathcal{H}=$ $\frac{1}{2\kappa}\frac{E_{i}^{a}E_{j}^{b}}{\sqrt{\text{det}\,h}}\epsilon^{ij}{}_{k}\left(\mathcal{F}_{ab}^{k}(A)-(1+\beta^{2})\epsilon^{k}{}_{mn}K_{a}^{m}K_{b}^{n}-2\frac{1+\beta^{2}}{\beta}\mathcal{D}_{[a}^{(\Gamma)}K_{b]}^{k}\right)$
$-\frac{i}{2}E_{i}^{a}\left(\overline{\Psi}\gamma^{i}D_{a}^{(A)}\Psi-\overline{D_{a}^{(A)}\Psi}\right)-\frac{1}{2}\epsilon^{i}{}_{jk}E_{i}^{a}K_{a}^{j}J^{k}+\frac{\beta}{2}E_{i}^{a}K_{a}^{i}J^{0}$ (50)
@@
In the above formulation, the canonical pair for the fermionic system is $(\Psi,\Pi)$ where $\Pi=i\sqrt{\text{det}\,h}\Psi$. However, this leads to a symplectic term in the action in which there is an additional term implying an imaginary correction to the new Ashtekar connection. This will ultimately lead to a modification of the Poisson brackets. This issue can be resolved by taking into account the half-density fermionic fields Bojowald:2010qpa ; Thiemann:1998QSDV ; Thiemann:1997rq . Thus, the new spinors and their conjugate momenta become
$\xi=$ $4]{\text{det}\,h}\Psi,$ $\pi=$ $i\xi^{\dagger}.$ (52)
@@
where $x,\,y$ are points residing on the same spatial hypersurface. Additionally, the fermionic axial 4-current can also be redefined by introducing the densitized axial 4-current
@@
Moreover, one can show that Thiemann:1998QSDV
@@
$ie_{0}^{a}\gamma^{0}D_{a}^{(A)}\Psi-\frac{1}{2}\beta e_{0}^{a}K_{a}^{i}\gamma_{5}\gamma_{i}\Psi+ie_{i}^{a}\gamma^{i}D_{a}^{(A)}\Psi+\frac{1}{2}\epsilon^{i}{}_{jk}e_{i}^{a}K_{a}^{j}\gamma_{5}\gamma^{k}\Psi-\frac{1}{2}\beta e_{i}^{a}K_{a}^{i}\gamma_{5}\gamma^{0}\Psi=0\,,$ (58)
@@
$(t^{a}-N^{a})[i\gamma^{0}D_{a}^{(A)}\xi-\frac{1}{2}\beta K_{a}^{i}\gamma_{5}\gamma_{i}\xi]-N[ie_{i}^{a}\gamma^{i}D_{a}^{(A)}\xi+\frac{1}{2}\epsilon^{i}{}_{jk}e_{i}^{a}K_{a}^{j}\gamma_{5}\gamma^{k}\xi-\frac{1}{2}\beta e_{i}^{a}K_{a}^{i}\gamma_{5}\gamma_{0}\xi]=0\,.$ (59)
@@
$\mathcal{H}=$ $\frac{1}{2\kappa}\frac{E_{i}^{a}E_{j}^{b}}{\sqrt{\text{det}\,h}}\epsilon^{ij}{}_{k}\left(\mathcal{F}_{ab}^{k}(A)-(1+\beta^{2})\epsilon^{k}{}_{mn}K_{a}^{m}K_{b}^{n}-2\frac{1+\beta^{2}}{\beta}\mathcal{D}_{[a}^{(\Gamma)}K_{b]}^{k}\right)$ (61)
$-\frac{1}{2}e_{i}^{a}\left(\pi\gamma^{0}\gamma^{i}D_{a}^{(A)}\xi-D_{a}^{(A)}\pi\gamma^{0}\gamma^{i}\xi\right)+\frac{i}{2}\epsilon^{i}{}_{jk}e_{i}^{a}K_{a}^{j}\pi\gamma^{0}\gamma^{k}\xi-\frac{i\beta}{2}e_{i}^{a}K_{a}^{i}\pi\gamma^{0}\gamma_{5}\gamma^{0}\xi\,,$
@@
$+\frac{1}{2}(-1+2\beta^{2})\frac{\kappa\beta^{2}}{1+\beta^{2}}\frac{1}{\sqrt{\text{det}\,h}}\left(\pi_{L}\tau_{i}\xi_{L}+\pi_{R}\tau_{i}\xi_{R}\right)\left(\pi_{L}\tau^{i}\xi_{L}+\pi_{R}\tau^{i}\xi_{R}\right)$
$+\beta e_{i}^{a}D_{a}^{(A)}\left(\pi_{L}\tau^{i}\xi_{L}+\pi_{R}\tau^{i}\xi_{R}\right)-\frac{\kappa}{2}(1+2\beta^{2})\mathcal{G}_{i}\left(\pi_{L}\tau^{i}\xi_{L}+\pi_{R}\tau^{i}\xi_{R}\right)$

### v2

# arXiv:2507.06276v2[gr-qc] 26 Feb 2026

@@ Abstract
We present the full Hamiltonian analysis of electrodynamics, including fermions, photons, and their interaction, on a general curved spacetime in Ashtekar-Barbero variables. Only the time gauge is imposed and the theory is background independent. The details of the constraints of the theory for fermionic and electromagnetic fields coupled to gravity are presented. Moreover, the equations of motion are derived and it is shown that they reduce to the Dirac equation, Maxwell equation and their interaction term in curved or flat spacetime in the appropriate limits. Finally, the behavior of the theory under parity transformation is analyzed. This result presents a suitable foundation for loop or polymer quantization.
@@ 1 Introduction
There are previous works done with regard to fermions in a canonical theory in Ashtekar-Barbero variables Bojowald (2010); Das (2008); Thiemann (1998b); Jacobson (1988); Mercuri (2006b, a); Date (2011); Romero et al. (2021); Duque (2025); Romero et al. (2022) that paved the way for our work. However, they usually ignore the interaction term between the fermions and the photons and hence do not present the full canonical electrodynamics (CED) Hamiltonian in a non-fixed general curved background. Some of them also lack the details of computation or seemingly suffer from some missing terms or some inconsistencies. In particular, it seems that in Das (2008) the choice of the definition of the spin connection $\Gamma_{a}^{i}$ is not compatible with their covariant derivative $D_{a}=\partial_{a}+A_{a}^{i}\tau_{i}$ and their constraints. In Thiemann (1998b), the authors present the Hamiltonian constraint for gravity plus fermions, however, it seems that they do not start directly from the action, and as a consequence there is a missing term in the modification of the Ashtekar connection as well as in the fermionic Hamiltonian constraint. They also seem to take the modification of the Ashtekar connection derived in Jacobson (1988). However, the latter starts from the Palatini action rather than the Holst action . As a consequence it does not seem to involve axial fermionic current and only incoporates fermionic current. It also seems that the fact that the extrinsic curvature in presence of fermions is no longer symmetric because of the torsion contribution is overlooked. As a consequence, in the scalar constraint, a term proportional to $\epsilon_{i}{}^{jk}K_{a}^{i}E_{j}^{a}J_{k}$ is missing.
@@
In this work we set to fill the aforementioned gaps. We carefully study a generic background-independent system of gravity, fermions and photons, starting from the action and making a $3+1$ decomposition, writing the theory in Ashtekar-Barbero variables. This is done for both Weyl and Dirac fermions, in a step by step, clear, and detailed manner. We then derive the modification to both the canonical variables, and the constraints, showing in a clear way the distinct contribution of each sector of to these constraints. We also provide the equations of motion of fermions in this setting and study the behavior of the system under parity transformation. Another difference of this work compared to the previous ones, is that we do not modify the gamma matrices, and rather change the identities between them to incorporate the changes needed for the background-independent framework. This way it is easier to go back to a flat field theory. Furthermore, as mentioned before, to our knowledge, this is the first time that the CED interaction term is studied in detail in such a canonical formulation in Ashtekar-Barbero variables.
@@
The structure of this paper is as follows: In Sec. 2, we provide a brief overview of vacuum canonical gravity in Ashtekar-Barbero variables starting from the Holst action. Then in Sec. 3, we present how fermions are incorporated in the theory, and how they modify the canonical variables and constraints. Using the results of the previous sections, in Sec. 4.1 the full CED Hamiltonian in an arbitrary dynamical background is constructed by incorporating both the photon field and its interaction with fermions and gravity. In Sec. 4.2 we study the behavior of the system under the parity transformations. Finally, in Sec. 5 we summarize and conclude the paper.
@@ 2 Brief overview of Ashtekar-Barbero variables
In this section we will quickly review Ashtekar-Barbero formalism of GR in order to introduce the notation used in the Fermions and CED Sections.
We start with a first order action of GR, the Holst action Bojowald (2010)
@@
Throughout this work the first four lower case Latin indices a,b,c,da,\,b,\,c,\,d are abstract spacetime (sometimes spatial) indices. The rest of the lower case Latin indices are $su(2)$ indices. Upper case Latin indices are the internal Lorentz indices associated to the $SO(1,3)$ gauge freedom under which the tetrads transform. The connection one form associated to this transformation is $\omega_{a}^{IJ}$, and $F^{IJ}(\omega)=d\omega^{IJ}+\omega^{I}{}_{K}\wedge\omega^{KJ}$ is the antisymmetric 2-curvature of $\omega_{a}^{IJ}$. Furthermore $\kappa=8\pi G$ and
@@
where the covariant derivative is defined as $\mathcal{D}_{a}v^{I}=\partial_{a}v^{I}+\omega_{a}^{I}\,{}_{J}v^{J}$, with the indices carrying their aforementioned meanings and $v^{I}$ being an $SU(2)$-vector. This condition expresses the compatibility between the tetrad and the spin connection. This ensures that all tensor fields (such as the Dirac spinor field, the metric, etc.) transform properly under local Lorentz transformations.
To proceed toward a canonical formulation, one foliates the spacetime into spatial slices following the ADM decomposition. For this purpose, a spacetime tensor field, the triad $\mathcal{E}_{I}^{a}=e_{I}^{a}+n_{I}n^{a}$, is introduced where $n^{a}$ is the unit timelike vector field normal to the spatial slices, and $n_{I}\coloneqq e_{I}^{a}n_{a}$. It is also common to fix the boost part of the internal Lorentz transformations by imposing the “time gauge” condition $e_{0}^{a}=n^{a}$ Bojowald (2010).
@@
where $N$ is the lapse function, $N^{a}$ the shift vector and $t^{a}$ is the time-evolution vector field defined such that $t^{a}\nabla_{a}t=1$. Making use of the relation $|\text{det}\,e|=N\sqrt{\text{det}\,h}$, where $\text{det}\,h$ is the determinant of the spatial metric $h_{ab}$ induced on the spatial hypersurfaces, the action becomes
@@
3 Fermionic Field
The final goal of the present manuscript is analyzing the CED framework (i.e., the case involving both fermions and photons and their interaction). All electrically charged fermions are described by Dirac spinors. In contrast, Majorana fermions —being their own antiparticles— must be electrically neutral. Since CED describes interactions between charged particles mediated by photon exchange, the theory inherently involves Dirac fermions. Consequently, we adopt the four-component representation of fermionic fields Pal (2011); Schwartz (2013). For further details, refer to Appendix A.
First of all, it is necessary to study how the inclusion of fermions alters the familiar vacuum scenario. This system has been previously studied in the literature Bojowald (2010); Das (2008); Thiemann (1998b); Jacobson (1988); Mercuri (2006b, a); Date (2011). Notably, in Bojowald (2010) and Das (2008), the Dirac matrices themselves were modified to accommodate the curved spacetime geometry. In contrast, the present work adopts an approach in which instead of altering the matrices, we modify the algebraic identities involving them (see Appendix A). This results in a formulation that more mathematically closely mirrors the flat spacetime treatment of fermionic fields, allowing a more straightforward comparison between the two cases.
To derive the constraints associated to the fermionic contribution we apply the canonical decomposition, $\mathcal{E}_{I}^{a}=e_{I}^{a}+n^{a}n_{I}$, to the Dirac action
where $n_{I}=-\delta_{I}^{0}$ (in the time gauge), as was done in the vacuum case. As a result, the fermionic action becomes
@@ 3 Fermionic Field
Once again we encounter three types of terms involving $t^{a},\,N$, and $N^{a}$ . The term containing $t^{a}$ leads to
@@
$\mathcal{H}^{F}=$ $-\frac{i}{2}E_{i}^{a}\left(\overline{\Psi}\gamma^{i}\mathfrak{D}_{a}\Psi-\overline{\mathfrak{D}_{a}\Psi}\gamma^{i}\Psi\right)+\sqrt{\text{det}\,h}\,m\,\overline{\Psi}\Psi\,$, (31)
@@
It is worth noticing that we recover the Weyl fermion case simply by setting $m=0$ in the above.
By minimally coupling the vacuum action (see Section 2) to the fermionic action, we obtain the total Hamiltonian of the system, comprising both the vacuum and the fermionic field contributions.
@@
For the sake of notational simplicity, we will omit the superscript “G+F” in the remaining of this section. Unless otherwise stated, all expressions should be understood as referring to the total system, including both the vacuum and the fermionic fields.
@@
This implies that the system is no longer torsionless and the canonical variables of the theory now carry an implicit dependence on the fermionic field. We emphasize that the present result is stated explicitly, with consistent conventions for both signatures and notation. In contrast, in Das (2008) we find some inconsistencies in the sign conventions, and the derivation of the result is not presented in a fully transparent way. Moreover, in Thiemann (1998b) the Hamiltonian constraint is computed for gravity coupled to fermions; however,it is not entirely clear to us whether this Hamiltonian is derived directly from the action. This may account for the absence of a term in the modification of the Ashtekar connection, as well as in the fermionic contribution to the Hamiltonian constraint. Moreover, as already noted in Perez and Rovelli (2006), unlike the vacuum case, where the Barbero-Immirzi parameter plays no role in the equations of motion, here the torsion term is proportional to this parameter, effectively governing the strength of the interaction between the gravitational and fermionic fields. As a result, to enforce the compatibility condition in the presence of fermions, the connection 1-form must be modified as
@@
$=\beta K_{a}^{i}+\tilde{\Gamma}_{a}^{i}+\frac{\kappa}{4}\frac{\beta^{2}}{1+\beta^{2}}\left(e_{a}^{i}J^{0}-\frac{1}{\beta}\epsilon^{i}{}_{jk}e_{a}^{j}J^{k}\right)\,.$
@@
Once again, we see that the $\mathcal{S}_{i}$ constraint (35) is proportional to the Gauss constraint and can be neglected. Finally, the first class constraints (detailed calculations in Appendix B) in the case in which gravity is minimally coupled to a fermionic field are
@@
[…]
@@
Another modification arises in the definition of the fermionic covariant derivatives, which can now be re-expressed in terms of the covariant derivative associated with the modified Ashtekar connection, along with an additional correction term
@@
As a side remark, we note that there appears to be some inconsistency in the literature regarding the definition of the fermionic covariant derivative. In particular, in Das (2008) the same definitions for the spin connection $\Gamma_{a}^{i}$ and the Pauli basis $\tau^{i}$ as those adopted here are used. However, when expressing the covariant derivative (49) in the Pauli basis, the authors obtain $D_{a}=\partial_{a}+A_{a}^{i}\tau_{i}$, which is not consistent with the stated definition of the covariant derivative, nor with the resulting expressions for the constraints, which otherwise agree with ours. A similar issue appears in Thiemann (1998b). In that case, however, the explicit definition of $\Gamma_{a}^{i}$ is not provided, preventing a direct comparison. We note that alternative sign conventions for the spin connection are sometimes adopted in the literature. For completeness, we have verified whether our results remain unchanged under such a modification. We find that not only do the resulting constraints differ from those obtained here, but also additional inconsistencies would arise already at the level of the vacuum constraints (e.g., the vacuum Gauss constraint does not take its broadly accepted form), which are generally assumed to be well established. Our understanding is that this discrepancy may originate from an ambiguity in index conventions, allowing for sign changes without immediate consequences in intermediate expressions. Since the explicit form of the fermionic covariant derivative in the Pauli basis is not required in the remainder of this work, we do not pursue this issue further and leave a more detailed investigation to future studies. Continuing the computations, for the conjugate fermionic field we find
@@
[…]
$-\frac{i}{2}E_{i}^{a}\left(\overline{\Psi}\gamma^{i}D_{a}^{(A)}\Psi-\overline{D_{a}^{(A)}\Psi}\right){+}\frac{1}{2}\epsilon^{i}{}_{jk}E_{i}^{a}K_{a}^{j}J^{k}+\frac{\beta}{2}E_{i}^{a}K_{a}^{i}J^{0}+\sqrt{\text{det}\,h}\,m\,\overline{\Psi}\Psi\,$, (51)
@@
In the above formulation, the canonical pair for the fermionic system is $(\Psi,\Pi)$ where $\Pi=i\sqrt{\text{det}\,h}\Psi$. However, this leads to a symplectic term in the action in which there is an additional term implying an imaginary correction to the new Ashtekar connection. This will ultimately lead to a modification of the Poisson brackets. This issue can be resolved by taking into account the half-density fermionic fields Bojowald (2010); Thiemann (1998b, a). Thus, the new spinors and their conjugate momenta become
$\xi=$ $\sqrt[4]{\text{det}\,h}\Psi,$ $\pi=$ $i\xi^{\dagger}.$ (53)
@@
where x,yx,\,y are points residing on the same spatial hypersurface. Additionally, the fermionic axial 4-current can also be redefined by introducing the densitized axial 4-current
@@
Moreover, one can show that Thiemann (1998b)
@@
$ie_{0}^{a}\gamma^{0}D_{a}^{(A)}\Psi-\frac{1}{2}\beta e_{0}^{a}K_{a}^{i}\gamma_{5}\gamma_{i}\Psi+ie_{i}^{a}\gamma^{i}D_{a}^{(A)}\Psi-\frac{1}{2}\epsilon^{i}{}_{jk}e_{i}^{a}K_{a}^{j}\gamma_{5}\gamma^{k}\Psi-\frac{1}{2}\beta e_{i}^{a}K_{a}^{i}\gamma_{5}\gamma^{0}\Psi=0\,$, (59)
@@
$(t^{a}-N^{a})[i\gamma^{0}D_{a}^{(A)}\xi-\frac{1}{2}\beta K_{a}^{i}\gamma_{5}\gamma_{i}\xi]-N[ie_{i}^{a}\gamma^{i}D_{a}^{(A)}\xi-\frac{1}{2}\epsilon^{i}{}_{jk}e_{i}^{a}K_{a}^{j}\gamma_{5}\gamma^{k}\xi-\frac{1}{2}\beta e_{i}^{a}K_{a}^{i}\gamma_{5}\gamma_{0}\xi]=0\,.$ (60)
@@
[…]
$-\frac{1}{2}e_{i}^{a}\left(\pi\gamma^{0}\gamma^{i}D_{a}^{(A)}\xi-D_{a}^{(A)}\pi\gamma^{0}\gamma^{i}\xi\right)-\frac{i}{2}\epsilon^{i}{}_{jk}e_{i}^{a}K_{a}^{j}\pi\gamma^{0}\gamma^{k}\xi-\frac{i\beta}{2}e_{i}^{a}K_{a}^{i}\pi\gamma^{0}\gamma_{5}\gamma^{0}\xi\,$,
@@
$+\frac{i}{2}\beta e_{i}^{a}D_{a}^{(A)}(\pi\gamma^{0}\gamma_{5}\gamma^{i}\xi)-\frac{1}{8}({3}+2\beta^{2})\frac{\kappa\beta^{2}}{1+\beta^{2}}\frac{1}{\sqrt{\text{det}\,h}}(\pi\gamma^{0}\gamma_{5}\gamma_{i}\xi)(\pi\gamma^{0}\gamma_{5}\gamma^{i}\xi)$
$-\frac{i\kappa}{4}({-3}+2\beta^{2})\mathcal{G}_{i}(\pi\gamma^{0}\gamma_{5}\gamma^{i}\xi)-\frac{1}{2}e_{i}^{a}\left(\pi\gamma^{0}\gamma^{i}D_{a}^{(A)}\xi-D_{a}^{(A)}\pi\gamma^{0}\gamma^{i}\xi\right)-im\pi\gamma^{0}\,$, (65)
In contrast to the massive case, when treating with Weyl fermions, it is convenient to use the bispinor notation and express the constraints in terms of the Pauli basis. i.e.
@@
$+\frac{1}{2}(3+2\beta^{2})\frac{\kappa\beta^{2}}{1+\beta^{2}}\frac{1}{\sqrt{\text{det}\,h}}\left(\pi_{L}\tau_{i}\xi_{L}+\pi_{R}\tau_{i}\xi_{R}\right)\left(\pi_{L}\tau^{i}\xi_{L}+\pi_{R}\tau^{i}\xi_{R}\right)$
$+\beta e_{i}^{a}D_{a}^{(A)}\left(\pi_{L}\tau^{i}\xi_{L}+\pi_{R}\tau^{i}\xi_{R}\right)-\frac{\kappa}{2}(-3+2\beta^{2})\mathcal{G}_{i}\left(\pi_{L}\tau^{i}\xi_{L}+\pi_{R}\tau^{i}\xi_{R}\right)$